# SparX: A Sparse Cross-Layer Connection Mechanism for Hierarchical Vision Mamba and Transformer Networks

### v1

@@ 4 Experiments
In this section, we conduct experimental evaluations on representative vision tasks, starting with image classification. Afterwards, the pre-trained models are transferred to downstream tasks, including object detection and semantic segmentation. All experiments were conducted on 8 NVIDIA H800 GPUs. Due to the page limit, we can only present part of the results in this section while more experimental results are provided in Appendix A.
@@ 4.1 Image Classification
Setup. For the image classification task, we use the ImageNet-1K dataset (Deng et al. 2009) and strictly follows the same experimental setting described in DeiT (Touvron et al. 2021) to ensure a fair comparison. Our method is compared with many representative models: CNN-based models (ConvNeXt (Liu et al. 2022), InceptionNeXt (Yu et al. 2024b), SLaK (Liu et al. 2023), and UniRepLKNet (Ding et al. 2024)), Transformer-based models (Focal-Transformer (Yang et al. 2021), PVTv2 (Wang et al. 2022), Swin (Liu et al. 2021), CSWin (Dong et al. 2022), and UniFormer (Li et al. 2023)), and Mamba-based models (PlainMamba (Yang et al. 2024), ViM2 (Behrouz, Santacatterina, and Zabih 2024), VMamba (Liu et al. 2024), EfficientVMamba (Pei, Huang, and Xu 2024), and LocalVMamba (Huang et al. 2024)).
Results. Table 2 demonstrates the significant advantages of our models over other CNN-, transformer-, and Mamba-based methods. For example, SparX-Mamba-T outperforms ConvNeXt-T/Swin-T by a large margin of 1.4%/2.2%, respectively, in top-1 accuracy. Compared to the recent VMamba-T, our model achieves a 1.0% higher top-1 accuracy with fewer Params and only a slight increase in FLOPs. Furthermore, even with larger models, SparX-Mamba maintains a clear advantage. It is worth noting that while our method’s complexity is aligned with VMamba, some representative methods included in Table 2 have clearly fewer FLOPs or Params. The performance of these methods is included in Table 2 for the sake of completeness, but not for direct comparison with our model. Meanwhile, we have also provided a more comprehensive comparison with more advanced CNN and transformer models (Yu et al. 2024a; Tu et al. 2022; Wang et al. 2023b; Zhu et al. 2023). As shown in Appendix A.5, our SparX-Mamba demonstrates competitive performance in ImageNet-1K classification and exhibits clear advantages on downstream tasks.
4.2 Object Detection and Instance Segmentation
Setup. To evaluate our network architecture on object detection and instance segmentation, we conduct experiments on the COCO 2017 dataset (Lin et al. 2014). We use the Mask R-CNN framework (He et al. 2017) and adopt the same experimental settings as in Swin (Liu et al. 2021). The backbone networks are initially pre-trained on ImageNet-1K and subsequently fine-tuned for 12 epochs (1$\times$ schedule) as well as 36 epochs (3$\times$ schedule + multi-scale training). Due to the page limit, results based on different IoU thresholds are given in Table 7 of Appendix.
@@ 4.2 Object Detection and Instance Segmentation
Table 3: Comparison of object detection and instance segmentation performance on the COCO dataset. FLOPs are calculated for the 800$\times$1280 resolution.
4.3 Semantic Segmentation
Setup. Experiments on semantic segmentation are conducted using the ADE20K dataset (Zhou et al. 2017). We employ two segmentation frameworks, Semantic FPN (S-FPN) (Kirillov et al. 2019) and UperNet (Xiao et al. 2018). To ensure fair comparisons, we initialize all backbone networks with ImageNet-1K pre-trained weights. Furthermore, we strictly adhere to the same training settings as outlined in previous work (Li et al. 2023).
Results. Due to the page limit, detailed experimental results on semantic segmentation are presented in Table 6 of the Appendix.
4.4 Ablation Studies
Setup. To assess the effectiveness of individual components in SparX, we conduct comprehensive ablation studies on image classification and semantic segmentation. We first train each model variant on the ImageNet-1K dataset using the same training setting as in Section 4.1. Then, we fine-tune the pre-trained models on the ADE20K dataset, using the S-FPN framework and maintaining an identical training configuration as described in Section 4.3. Due to the page limit, we only present part of ablation studies here, additional results can be found in Appendix A.4.
Method ImageNet-1K ADE20K
Mem T F P Acc. mIoU
(MB) (imgs/s) (G) (M)
VMamba-T 6784 1613 4.9 31.0 82.5 47.2
DGC-Mamba-T 7523 860 5.5 28.8 83.4 49.3
DSN-Mamba-T 7641 830 6.0 30.9 83.4 49.5
SparX-Mamba-T 7066 1370 5.2 27.1 83.5 49.5
Table 4: Comparision with DenseNet-like networks. The throughput (T) and memory (Mem) usage are evaluated on a single H800 GPU using a batch size of 128. Due to space constraints, we omit the complexity of the segmentation network. However, it is noteworthy that its complexity is closely aligned with that of the classification network.
Method ImageNet-1K ADE20K
Mem T F P Acc. mIoU
(MB) (imgs/s) (G) (M)
Concat 6861 1413 5.5 27.9 82.8 48.3
w/o GCCA 6853 1425 5.3 26.4 82.7 48.1
SCA 12453 1275 5.2 27.9 83.3 49.2
SRA 9333 997 5.3 27.0 83.2 49.2
DMCA 7066 1370 5.2 27.1 83.5 49.5
Table 5: Comparision with different cross-layer interaction methods.
Comparsion with DenseNet-like networks. Based on SparX-Mamba-T, we conduct the following experiments: (1) To validate our sparse cross-layer connections, we maintain the stride (S) = 2 in SparX, and set up cross-layer connections by removing the sliding window and connecting each ganglion layer with all preceding ganglion and normal layers. This design is denoted as “Dense Ganglion Connections (DGC-Mamba-T)”. (2) To verify the superiority of our model over DenseNet (Huang et al. 2017), we set S=1 and remove the cross-layer sliding window, thereby designing a network similar to DenseNet. This model is referred to as “DenseNet-style Network (DSN-Mamba-T)”.
Comparsion with different interaction methods. Based on SparX-Mamba-T, we replace our DMCA with other fusion methods for multi-layer feature interactions: (1) Remove the entire DMCA module and directly concatenate features from the current layer and selected preceding layers, followed by a Conv-FFN layer (Wang et al. 2022) for feature fusion. Using Conv-FFN for feature fusion maintains a comparable level of complexity with the baseline model. This design is referred to as “Concat”. (2) Remove the grouped channel-wise cross-attention (GCCA) operation in DMCA. That means there is no need to perform spatial reduction and compute the cross-attention matrix between spatially reduced query and key, but value still needs to be computed. Then query and value are directly concatenated and fused using a Conv-FFN layer. This design is denoted as “w/o GCCA”. (3) Replace the attention computation in DMCA with Spatial Reduction Attention (SRA) proposed in PVT (Wang et al. 2021) to compare the performance of channel mixing and spatial mixing. (4) Replace our DMCA with spatial cross-attention (SCA) proposed in FcaFormer (Zhang, Hu, and Wang 2023) to compare the performance of different multi-layer feature aggregation mechanisms. This version is denoted as “SCA”.
As listed in Table 5, our DMCA demonstrates better cross-layer feature aggregation and selection thanks to dynamic feature retrieval from preceding layers. Specifically, both “Concat” and “w/o GCCA” result in obvious performance degradation compared with our DMCA. Meanwhile, DMCA outperforms both SRA and SCA possibly because DMCA offers dynamic channel mixing, which is complementary to spatial mixing performed with SSMs. It is noteworthy that some previous works have demonstrated that incorporating both channel and spatial dynamics can lead to improved performance (Fu et al. 2019; Ding et al. 2022a). Conversely, both SRA and SCA function as a spatial mixer similar to SSM, compromising the model’s ability to effectively represent multidimensional features. More importantly, compared with simple feature concatenation, both SRA and SCA incur higher memory costs and lower throughput due to quadratic complexity, whereas our DMCA only marginally increases computational overhead, indicating greater efficiency.
5 Conclusion
In this work, we propose a new skip-connection strategy named SparX, drawing inspiration from the Retinal Ganglion Cell (RGC) layer in the human visual system. SparX aims to create sparse cross-layer connections to enhance information flow and promote feature distillation and reuse in Mamba-based models. In addition, we propose a Dynamic Multi-layer Channel Aggregator (DMCA) that facilitates dynamic feature aggregation and interaction across layers. On the basis of SparX, we further propose SparX-Mamba and SparX-Swin network architectures, showcasing superior performance on various challenging vision tasks.
Appendix A Appendix
A.1 Performance of Semantic Segmentation
@@
Table 6: Comparison of semantic segmentation performance on the ADE20K dataset. FLOPs are calculated for the 512$\times$2048 resolution. $\dagger$: baselines implemented by ourselves. For UperNet, we report both single-scale and mulit-scale mIoU.
@@
As shown in Table 6, our SparX-Mamba demonstrates significant advantages when integrated into both S-FPN and UperNet. For instance, when UperNet is used, SparX-Mamba increases mIoU by 1.7%/2.2% over VMamba-T. Performance improvements achieved with small/base models are also evident. Meanwhile, when using S-FPN, SparX-Mamba-T improves VMamba-T by a notable 2.3% mIoU and outperforms CSWin-T by 1.3% mIoU. These advantages are also evident in small and base models, effectively demonstrating the robustness of our model in dense prediction tasks.
A.2 More Detailed Comparisons in Object Detection and Instance Segmentation
A.3 Versatility Analysis
@@
Setup. To demonstrate the versatility of our cross-layer connectivity pattern, we have conducted experiments by incorporating SparX into networks with different token mixer architectures, including self-attention, vanilla SSM in Mamba, and Bi-SSM in ViM. To ensure a fair comparison with other hierarchical networks, we follow the conventional design, where a token mixer is followed by an FFN, with the token mixer having the aforementioned multiple options. Specifically, taking SparX-Mamba-T as a basic network, we replace the SS2D block with shifted window attention proposed in Swin Transformer (Liu et al. 2021) to obtain a network denoted as SparX-Swin-T (i.e., Shifted Window Attention $\rightarrow$ FFN), and increase the number of channels in stage 4 from 512 to 576 to align its computational cost with Swin-T. Next, we scale up the model size of SparX-Swin-T by increasing the depths to [2, 2, 27, 3] and channels to [120, 240, 396, 616] to obtain a larger model, denoted as SparX-Swin-B, which has a model complexity comparable to that of Swin-B. Moreover, we also replace the SS2D module in VSS with the vanilla SSM module in original Mamba (Gu and Dao 2024) (i.e., SSM $\rightarrow$ FFN) and Bi-SSM module in ViM (Zhu et al. 2024) (i.e., Bi-SSM $\rightarrow$ FFN). For fair comparisons, we have strictly maintained the similarity of other micro-blocks, such as removing DPE and using identical patch embedding layers. We first train each model variant on the ImageNet-1K dataset, and then fine-tune the pre-trained models on the ADE20K dataset using the UperNet framework.

### v2

@@ 4 Experiments
In this section, we present comprehensive experimental evaluations on representative vision tasks, starting with image classification. Afterwards, the pre-trained models are transferred to downstream tasks, including object detection and semantic segmentation. All experiments were conducted on 8 NVIDIA H800 GPUs. More experimental results are provided in Appendix A.
@@ 4.1 Image Classification
Setup. We use the ImageNet-1K dataset (Deng et al. 2009) and follow the same experimental setting described in DeiT (Touvron et al. 2021) for a fair comparison. Our method is compared with many representative vision models: CNN-based models (ConvNeXt (Liu et al. 2022), InceptionNeXt (Yu et al. 2024b), SLaK (Liu et al. 2023), UniRepLKNet (Ding et al. 2024), and MambaOut (Yu and Wang 2024)), Transformer-based models (Focal-Transformer (Yang et al. 2021), PVTv2 (Wang et al. 2022), Swin (Liu et al. 2021), CSWin (Dong et al. 2022), and UniFormer (Li et al. 2023)), and Mamba-based models (PlainMamba (Yang et al. 2024), ViM2 (Behrouz, Santacatterina, and Zabih 2024), VMamba (Liu et al. 2024), EfficientVMamba (Pei, Huang, and Xu 2024), LocalVMamba (Huang et al. 2024), and MSVMamba (Shi, Dong, and Xu 2024)).
Results. Table 2 demonstrates the significant advantages of our models over other CNN-, transformer-, and Mamba-based methods. For example, SparX-Mamba-T outperforms ConvNeXt-T/Swin-T by a large margin of 1.4%/2.2%, respectively, in top-1 accuracy. Compared to the recent VMamba-T, our model achieves a 1.0% higher top-1 accuracy with fewer Params and only a slight increase in FLOPs. Furthermore, even with larger models, SparX-Mamba maintains a clear advantage. Meanwhile, we have also provided a more comprehensive comparison with more advanced vision models in Appendix A. In general, our method demonstrates competitive performance in ImageNet-1K classification and exhibits clear advantages on downstream tasks.
@@
Table 3: Comparison of object detection and instance segmentation performance on the COCO dataset using Mask R-CNN framework. FLOPs are calculated for the 800$\times$1280 resolution.
4.2 Object Detection and Instance Segmentation
Setup. To evaluate our network architecture on object detection and instance segmentation, we conduct experiments on the COCO 2017 dataset (Lin et al. 2014). We use the Mask R-CNN framework (He et al. 2017) and adopt the same experimental settings as in Swin (Liu et al. 2021). The backbone networks are initially pre-trained on ImageNet-1K and subsequently fine-tuned for 12 epochs (1$\times$ schedule) as well as 36 epochs (3$\times$ schedule + multi-scale training).
@@ 4.2 Object Detection and Instance Segmentation
Table 4: Comparison of semantic segmentation performance on the ADE20K dataset. FLOPs are calculated for the 512$\times$2048 resolution. $\dagger$: baselines implemented by ourselves. For UperNet, we report both single-scale and multi-scale mIoU.
4.3 Semantic Segmentation
Setup. Experiments on semantic segmentation are conducted using the ADE20K dataset (Zhou et al. 2017). We employ two segmentation frameworks, Semantic FPN (S-FPN) (Kirillov et al. 2019) and UperNet (Xiao et al. 2018). To ensure fair comparisons, we initialize all backbone networks with ImageNet-1K pre-trained weights. Furthermore, we strictly adhere to the same training settings as outlined in previous work (Liu et al. 2021; Li et al. 2023).
@@ 4.3 Semantic Segmentation
Results. As shown in Table 4, our SparX-Mamba demonstrates significant advantages when integrated into both S-FPN and UperNet. For instance, when UperNet is used, SparX-Mamba increases mIoU by 1.7%/2.2% over VMamba-T. Performance improvements achieved with small/base models are also evident. Meanwhile, when using S-FPN, SparX-Mamba-T improves VMamba-T by a notable 2.3% mIoU and outperforms CSWin-T by 1.3% mIoU. These advantages are also evident in small and base models, effectively demonstrating the robustness of our model in dense prediction tasks.
4.4 Ablation Studies
Setup. To assess the effectiveness of individual components in SparX, we conduct comprehensive ablation studies on image classification and semantic segmentation. We first train each model variant on the ImageNet-1K dataset using the same training setting as in Section 4.1. Then, we fine-tune the pre-trained models on the ADE20K dataset, using the S-FPN framework and maintaining an identical training configuration as described in Section 4.3. We have also provided more analytical experiments, which can be found in Appendix A.
Method ImageNet-1K ADE20K
Mem T F P Acc. mIoU
(MB) (imgs/s) (G) (M)
VMamba-T 6784 1613 4.9 31.0 82.5 47.2
DGC-Mamba-T 7523 860 5.5 28.8 83.4 49.3
DSN-Mamba-T 7641 830 6.0 30.9 83.4 49.5
SparX-Mamba-T 7066 1370 5.2 27.1 83.5 49.5
Table 5: Comparison with DenseNet-like networks. The throughput (T) and memory (Mem) usage are evaluated on a single H800 GPU using a batch size of 128. Due to space constraints, we omit the complexity of the segmentation network. However, it is noteworthy that its complexity is closely aligned with that of the classification network.
Comparison with DenseNet-like networks. Based on SparX-Mamba-T, we conduct the following experiments: (1) To validate our sparse cross-layer connections, we maintain the stride (S) = 2 in SparX, and set up cross-layer connections by removing the sliding window and connecting each ganglion layer with all preceding ganglion and normal layers. This design is denoted as “Dense Ganglion Connections (DGC-Mamba-T)”. (2) To verify the superiority of our model over DenseNet (Huang et al. 2017), we set S=1 and remove the cross-layer sliding window, thereby designing a network similar to DenseNet. This model is referred to as “DenseNet-style Network (DSN-Mamba-T)”.
Method ImageNet-1K ADE20K
Mem T F P Acc. mIoU
(MB) (imgs/s) (G) (M)
Concat 6861 1413 5.5 27.9 82.8 48.3
w/o CGCA 6853 1425 5.3 26.4 82.7 48.1
CSA 7053 600 5.2 26.0 83.1 49.0
SRA 9333 997 5.3 27.0 83.2 49.2
SCA 12453 1275 5.2 27.9 83.3 49.2
DMCA 7066 1370 5.2 27.1 83.5 49.5
Table 6: Comparison with different cross-layer interaction methods.
Comparison with different interaction methods. Based on SparX-Mamba-T, we replace our DMCA with other fusion methods for multi-layer feature interactions: (1) Remove the entire DMCA module and directly concatenate features from the current layer and selected preceding layers, followed by a Conv-FFN layer (Wang et al. 2022) for feature fusion. Using Conv-FFN for feature fusion maintains a comparable level of complexity with the baseline model. This design is referred to as “Concat”. (2) Remove the cross-layer grouped channel attention (CGCA) operation in DMCA. That means there is no need to perform spatial reduction and compute the cross-attention matrix between spatially reduced query and key, but value still needs to be computed. Then query and value are directly concatenated and fused using a Conv-FFN layer. This design is denoted as “w/o CGCA”. (3) Replace the attention computation in DMCA with Spatial Reduction Attention (SRA) (Wang et al. 2021, 2022) and Cross-Scale Attention (CSA) (Shang et al. 2023) to compare the performance of channel mixing and spatial mixing. (4) Replace our DMCA with spatial cross-attention (SCA) proposed in FcaFormer (Zhang, Hu, and Wang 2023) to compare the performance of different multi-layer feature aggregation mechanisms. This version is denoted as “SCA”.
As listed in Table 6, our DMCA demonstrates better cross-layer feature aggregation and selection thanks to dynamic feature retrieval from preceding layers. Specifically, both “Concat” and “w/o CGCA” result in obvious performance degradation compared with our DMCA. Meanwhile, DMCA outperforms both SRA and SCA possibly because DMCA offers dynamic channel mixing, which is complementary to spatial mixing performed with SSMs. It is noteworthy that some previous works have demonstrated that incorporating both channel and spatial dynamics can lead to improved performance (Fu et al. 2019; Ding et al. 2022a). Conversely, both SRA and SCA function as a spatial mixer similar to SSM, compromising the model’s ability to effectively represent multidimensional features. More importantly, compared with simple feature concatenation, both SRA and SCA incur higher memory costs and lower throughput due to quadratic complexity, whereas our DMCA only marginally increases computational overhead, indicating greater efficiency.
Method ImageNet-1K ADE20K
F (G) P (M) Acc. F (G) P (M) mIoU
Baseline 5.2 27.1 83.5 197 30.8 49.5
PlainNet 5.4 27.4 82.4(-1.1) 199 31.1 47.7(-1.8)
S=1 5.7 29.9 83.5(+0.0) 206 33.7 49.7(+0.2)
S=3 5.1 26.3 83.2(-0.3) 194 30.1 49.0(-0.5)
S=4 5.1 25.8 82.9(-0.6) 193 29.6 48.7(-0.8)
M=2 5.2 26.8 83.0(-0.5) 197 30.6 48.6(-0.9)
M=4 5.5 27.5 83.5(+0.0) 198 31.2 49.7(+0.2)
Table 7: Ablation study on hyperparameters of the SparX.
Method ImageNet-1K ADE20K
F (G) P (M) Acc. F (G) P (M) mIoU
Baseline 5.2 27.1 83.5 197 30.8 49.5
w/o DPE 5.2 27.0 83.3 (-0.2) 196 30.7 49.0 (-0.5)
w/o Skip 5.2 27.0 83.1 (-0.4) 200 30.8 48.9 (-0.6)
w/o SR 5.4 27.0 83.4 (-0.1) 196 30.8 49.3 (-0.2)
Table 8: Ablation study on DMCA and DPE.
Hyperparameter settings in SparX. We conduct an in-depth exploration of hyperparameter settings in the proposed sparse cross-layer connection mechanism. By using the tiny model as the baseline (i.e., S=2, M=3), we conceive the following model variants: (1) All ganglion layers are replaced with normal layers, i.e., the network no longer has any cross-layer connections. This design is denoted as “PlainNet”. For a fair comparison, we make models considered in this comparison have comparable complexity by adjusting the number of channels in stage 3 and stage 4 of PlainNet to 384 and 576, respectively. (2) Keeping M=3 fixed, we vary the stride hyperparameter (S) controlling the sparsity of ganglion layers to 1, 3, and 4, respectively, to assess the impact of stride on final performance. (3) Keeping S=2 fixed, we investigate the impact of the cross-layer sliding window size by setting M to 2 and 4, respectively.
Results. As shown in Table 7, having dense ganglion layers (i.e., S=1) does not result in an obvious performance improvement even though the model complexity increases. This indicates that our sparse placement of ganglion layers already facilitates cross-layer communication. However, as S further increases, model performance gradually declines, suggesting that larger gaps between ganglion layers adversely affects the performance. Note that although S=3 is the second-best option, considering the increased complexity of deeper models, we set S=3 in our small and base versions for higher computational efficiency. Furthermore, experiments demonstrate that the optimal number of preceding ganglion layers in the cross-layer sliding window is 3, as further increasing M to 4 does not yield any obvious performance gains.
Method ImageNet-1K UperNet 160K Mask R-CNN
F (G) P (M) Acc. F (G) P (M) mIoU F (G) P (M) $AP_{1\times}^{b}$ $AP_{1\times}^{m}$ $AP_{3\times}^{b}$ $AP_{3\times}^{m}$
Swin-T 4.5 29 81.3 945 60 44.5 48 267 42.7 39.3 46.0 41.6
SparX-Swin-T 4.7 26 82.6(+1.3) 948 57 45.5(+1.0) 45 270 44.1(+1.4) 40.8(+1.5) 47.6(+1.6) 43.0(+1.4)
Swin-B 15.4 88 83.5 1188 121 48.1 107 496 46.9 42.3 48.6 43.3
SparX-Swin-B 16.2 88 83.9(+0.4) 1201 118 48.6(+0.5) 107 518 48.1(+1.2) 42.9(+0.6) 49.1(+0.5) 43.7(+0.4)
Table 9: Performance Comparison when SparX is applied to Vision Transformers.
Ablations of DMCA and DPE. By using the tiny model as the baseline, we further investigate the effect of DMCA and DPE (Chu et al. 2022) by evaluating the performance of the following alternative model designs. (1) Remove the DPE module, and the resulting model is denoted as “w/o DPE”. (2) Investigate the necessity of concatenating the output of GCCA ($\mathbf{Z}$) with query and value by only using the output of GCCA without concatenating it with X and $\mathbf{Y_{v}}$. This design is denoted as “w/o Skip”. (3) Eliminate the spatial reduction (SR) operations, and use the original spatial resolution of query and key when computing GCCA. This design is denoted as “w/o SR”. According to the results given in Table 8, the use of DPE boosted the top-1 accuracy by 0.2% and mIoU by 0.5%. Then, we can find that “w/o Skip” also gives rise to a performance drop, reflecting the importance of the shortcut connection. Moreover, although spatial reduction (SR) compresses spatial tokens, it actually slightly improves performance. One possible explanation is that utilizing compressed information is sufficient when estimating image-level feature similarities.
Method ImageNet-1K UperNet 160K
F (G) P (M) Acc. F (G) P (M) mIoU
VMamba-T
(SS2D) 4.9 31 82.5 948 62 48.3
SparX-Mamba
-T (SSM) 4.8 25 82.9(+0.4) 946 55 48.1(-0.2)
SparX-Mamba
-T (Bi-SSM) 5.1 25 83.1(+0.6) 951 55 48.7(+0.4)
SparX-Mamba
-T (SS2D) 5.2 27 83.5(+1.0) 954 57 50.0(+1.7)
4.5 Versatility Analysis
@@
Setup. To demonstrate the versatility of our cross-layer connectivity pattern, we have conducted experiments by incorporating SparX into networks with different token mixer architectures, including self-attention, vanilla SSM in Mamba, and Bi-SSM in ViM. To ensure a fair comparison with other hierarchical networks, we follow the conventional design, where a token mixer is followed by an FFN, with the token mixer having the aforementioned multiple options. Specifically, taking SparX-Mamba-T as a basic network, we replace the SS2D block with shifted window attention proposed in Swin Transformer (Liu et al. 2021) to obtain a network denoted as SparX-Swin-T (i.e., Shifted Window Attention $\rightarrow$ FFN), and increase the number of channels in stage 4 from 512 to 576 to align its computational cost with Swin-T. Next, we scale up the model size of SparX-Swin-T by increasing the depths to [2, 2, 27, 3] and channels to [120, 240, 396, 616] to obtain a larger model, denoted as SparX-Swin-B, which has a model complexity comparable to that of Swin-B. Moreover, we also replace the SS2D module in VSS with the vanilla SSM module in original Mamba (Gu and Dao 2024) (i.e., SSM $\rightarrow$ FFN) and Bi-SSM module in ViM (Zhu et al. 2024) (i.e., Bi-SSM $\rightarrow$ FFN). For fair comparisons, we have strictly maintained the similarity of other micro-blocks, such as removing DPE and using identical patch embedding layers.
Results. As shown in Table 9, SparX-Swin-T/B models outperform their vanilla counterparts on both image classification and dense prediction tasks while maintaining similar computational costs. This demonstrates that our method is general and equally applicable to Transformer-based models. Meanwhile, as shown in Table 10, although vanilla SSM and Bi-SSM do not utilize the DWConv employed in SS2D of VMamba, our SparX enables both vanilla SSM and Bi-SSM to outperform VMamba-T in terms of accuracy and computational efficiency. On the semantic segmentation task, the performance of SparX-Mamba-T (SSM) marginally lags behind VMamba-T. This is because causal modeling in vanilla SSM fails to capture sufficient contextual information. When the Bi-SSM module in ViM is used instead, the mIoU improves significantly by 0.4%, demonstrating better performance than VMamba-T.
5 Conclusion
In this work, we propose a new skip-connection strategy named SparX, drawing inspiration from the Retinal Ganglion Cell (RGC) layer in the human visual system. SparX aims to create sparse cross-layer connections to enhance information flow and promote feature distillation and reuse in vision backbone networks. In addition, we propose a Dynamic Multi-layer Channel Aggregator (DMCA) that facilitates dynamic feature aggregation and interaction across layers. Based on SparX, we further propose SparX-Mamba and SparX-Swin, both of which demonstrate superior performance across a range of challenging vision tasks.
Appendix A Appendix